E/D Night - UT 27-October-2020

1st half LBTO/Eng

2nd half LBTI Science

Night Center 00:004 - handover to LBTI @23:59 (10 min instrument switch, split the difference)



Night Plan

Rollover Plans from Previous night
  • AO on-sky work
    • OSCO at twilight (0.5hr)
    • Uranus Observatons (1hr)
      • /home/lbto/Service_Observing/Eng_D_Oct_25_31_2020/Mars_Uranus_Scripts
    • GRB 200826A - H-band Imaging with Bin 4 AO - (2hr)
      • /home/lbto/Service_Observing/Eng_D_Oct_25_31_2020/GRB_200826A
    • Mars Observations (1hr)
      • /home/lbto/Service_Observing/Eng_D_Oct_25_31_2020/Mars_Uranus_Scripts/
      • 2020-Oct-27 05:45 - 06:45 UT -local: 22:45-23:45
      • Plan for Observing:
        • Scripts manually edited to allow for H2+ND2
        • NSIGUI to Phobos (this is the AO Ref star)
        • Close the loop on Phobos
        • Use Telsvc for LUCI to offset to Mars ~15" away, absorb the offset
        • Continue script
    • ESM tests (time permitting) - can be executed in poor seeing and/or clouds
      • Located in: /home/lbto/Service_Observing/Eng_D_Oct_25_31_2020/ESM (OBS-N)
      • M52 (sets 2nd half) AORef=10.7 and 15.2
      • M34 (rises above elevation 19:30hr) AORef = 12.4
  • Poor Weather Backup
    • Science Backup
      • /home/lbto/Service_Observing/Eng_D_Oct_25_31_2020/Poor_Weather
      • H-band Library (LUCI)
      • ICRF3 Followup (LUCI)
  • Closed Dome
    • LUCI SOUL NCPA Measurements
    • LUCI2 N30 calibration updates

Night Log

LUCI SOUL Checks with ARGOS Source
  • Unable to find the Argos source
  • Extensive troubleshooting: Tried older Hexapos poisitons, homing filters, tried to look for source on LUCI N375 and it was not visible. Gustavo restarted some things, but verified that communications looked good and we should see it. Steve went up verify the source visibly but no source was seen. Suspicion is lamp is burnt out. IT8226. Gustavo will head up tomorrow
  • LUCI2 N30 Flats. I was able to determine times for missing filters before condensation built up on the window. We have been sitting at 100% humidity so not too surprising. The library has been updated and posted (and tested this time)

-- BarryRothberg - 26 Oct 2020
Topic revision: r4 - 27 Oct 2020, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback